# Versatile photonic frequency synthetic dimensions using a single Mach-Zehnder-interferometer-assisted device on thin-film lithium niobate

### v1

@@ Abstract
Investigating physical models with photonic synthetic dimensions has been generating great interest in vast fields of science. The rapid developing thin-film lithium niobate (TFLN) platform, for its numerous advantages including high electro-optic coefficient and scalability, is well compatible with the realization of synthetic dimensions in the frequency together with spatial domain. While coupling resonators with fixed beam splitters is a common experimental approach, it often lacks tunability and limits coupling between adjacent lattices to sites occupying the same frequency domain positions. Here, on the contrary, we conceive the resonator arrays connected by electro-optic tunable Mach-Zehnder interferometers in our configuration instead of fixed beam splitters. By applying bias voltage and RF modulation on the interferometers, our design extends such coupling to long-range scenario and allows for continuous tuning on each coupling strength and synthetic effective magnetic flux. Therefore, our design enriches controllable coupling types that are essential for building programmable lattice networks and significantly increases versatility. As the example, we experimentally fabricate a two-resonator prototype on the TFLN platform, and on this single chip we realize well-known models including tight-binding lattices, topological Hall ladder and Creutz ladder. We directly observe the band structures in the quasi-momentum space and important phenomena such as spin-momentum locking and the Aharonov-Bohm cage effect. These results demonstrate the potential for convenient simulations of more complex models in our configuration.
@@ Introduction
Fig. 1: Schematic and experimental setup. (A) Configuration of a lattice network in frequency synthetic dimensions. The lattice network is simulated by a train of modulated resonators. Instead of fixed beam splitters (BS), MZIs tunable by DC and RF signals are used to couple the adjacent resonators (purple boxes). The lattice network consequently contains three types of coupling stemming from local modulation on the resonators (orange), DC (cyan) and RF (green) signals applied on the MZIs, respectively. (B) Schematic of the Creutz ladder. The coupling types are of same definition in A and $\phi_{1}$, $\phi_{2}$ and $\phi_{3}$ are the potential gauges. (C) Experimental setup for realizing B with two MZI assisted race-track resonators on TFLN. Three sets of electrical signals consisting of DC and RF parts are applied on three sets of the ground-signal-ground (GSG) arranged electrodes through Bias-Tees to control the two resonators ($S_{1},S_{3}$) and the MZI ($S_{2}$). For detecting band structures, a probe laser with detuning $\Delta\omega$ is injected to excite the bands. The time- (quasi-momentum-) varying transmittance signal, which is a slice of the band structure, is detected by a photodetector (PD) followed by an oscilloscope triggered by the RF source. For obtaining mode distribution, the measurement part is replaced by scanning a Fabry-Perot (F-P) cavity together with a photodector.
@@
Creating a controllable artificial system, of which the dynamics highly resembles the one in real space, is a commonly used method to simulate inaccessible, complex and high-dimensional systems. The artificial systems also called synthetic dimensions, are therefore capable to investigate fascinating novel physics in the field of such as solid-state physics [1, 2, 3, 4]. In terms of realization, photonic system hold advantages since it has multiple manipulable degrees of freedom including polarization, space, frequency and orbital angular momentum [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Configuring synthetic dimensions in the form of optical frequency lattices is one of the popular approaches in pioneering works, where optical-fiber-based resonators are mainly exploited [17, 18, 19, 20, 21, 16, 22].
@@
Apart from optical fibers, integrated optics holds promise for generating even more complex and programmable lattice networks for its good scalability, stability and reconfigurablity. Especially, with broad transparent spectrum, high non-linear and electro-optic coefficients, thin-film lithium niobate (TFLN) provides an ideal platform. The rapid development of high performance devices on such platform including modulated resonators have paved the way for frequency synthetic dimensions [23, 24, 25, 26, 27, 28, 29, 30, 31]. As a result, constructing frequency lattices on TFLN has garnered increasing interest recently and several experimental works, such as Ref. [32, 33, 34, 35, 36] and our previous work Ref. [37], have been reported.
To extend the dimensions of lattices and interaction in synthetic dimensions, the realization of coherent coupling between resonators is crucial. When two resonators are connected with a passive coupler, adjacent coupling emerges between two lattices (resonators) at the same-frequency lattices points. It allows for the combination of the frequency and spatial dimension for simulating various models and engrossing phenomena, e.g., the Hall ladder, honeycomb lattices, band structures with effective magnetic field, and spin-momentum locking, etc. [18, 31]. Moreover, controllable long-range coupling, such as cross coupling, is crucial for more general models. For example, the historically important model of Creutz ladder [38, 39, 40, 41] requires the cross coupling between different-frequency points in different lattices, but still awaits experimental investigation leveraging frequency synthetic dimensions since the lack of sufficiently tunable coupling methods. Therefore, improving the tunabilitiy of the coupling modes between resonators is necessary to tune the lattice and realize more physical model.
@@
In this work, we design and fabricate a Mach-Zehnder-interferometer (MZI)-assisted double-resonator device on TFLN, and then realize the sufficient regulation of coupling strength and range in synthetic dimensions. By tuning the bias voltage applied on the MZI [42], we are able to arbitrarily tune the coupling strength between two modulated resonators and explore every model mentioned above with this single device. More importantly, by modulating the MZI rather than resonators with RF signals, we further extend the adjacent coupling to cross coupling and long-range coupling, without introducing auxiliary resonators [43, 44]. This enable us to simulate more complex modle like Creutz ladder. We show the experimental observation of band structures of tight-binding lattices, the Hall ladder and the Creutz ladder, as well as topological features including gauge potential and Aharonov-Bohm cage. While maintaining a similar scale and fabrication difficulty to using fixed beam splitters as couplers, our configuration significantly increases the versatility of devices for realizing frequency synthetic dimensions and broadens the scope of models that can be simulated.
@@
Fig. 2: Experimental obtained band structures of tight-binding lattices and the Hall ladder. The band structures exhibit tight-binding lattice characters when two resonators are either not coupled (A) or fully coupled (B). $J_{B}^{H}$ is set to zero in the left plot of (B), while being equal to $J_{A}^{H}$ in the right plot which consequently doubles the coupling strength. (C) The upper and the lower plots are respectively the numerical calculations of (A) and the left plot of (B), using a photonic integrated circuit design software. (D) Illustration of the Hall ladder where the crossing coupling is absent. (E to H) The heat maps are the experimentally obtained band structures of the Hall ladder with coupling strength ratio $r_{VH}=J_{V}/2J_{H}=0.75$, $J^{H}=0.06\Omega$ and an effective magnetic flux $\phi_{1}$ being ±$\pm\pi/4$ or ±$\pm 3\pi/4$. The lines on them are the corresponding theoretical bands solved from the $k$-space Hamiltonian. The distribution information on the two spins (resonators) are encoded in the normalized intensities of both the experimental and theoretical normalized intensities, represented by two sets of colors. Choosing opposite fluxes is equivalent to selectively exciting the opposite spin (injecting into the other resonator). For one spin (E and F), the negative (positive) $k$ states in the first Brillouin zone predominate the upper (lower) bands. Meanwhile, complementary patterns appear in the other spin (G and H), signifying the spin-momentum locking. (I and J) Using $\phi_{1}=\pi/2$ as an example, $r_{VH}$ is increase from $0.75$ to $1.46$ and the spin-momentum locking effect is reduced.
@@
Design of the couplings among the lattice network points is one of the core ingredients of the configuration of physical models, specifically, both the coupled modes and coupling range are required to be tuned appropriately. Towards an ideal versatile and programmable network, a device being capable to cover as many coupling types as possible is on demand. Now we consider a 2D network protocol constituted by a series of identical adjacently-coupled 1D lattices, which can be simulated by a train of resonators (Fig. 1A), by integrating the frequency and the spatial domain. We can depict the general model with the horizontal, vertical, crossing coupling strengths ($J^{H}_{i,p}$, $J^{V}_{i}$, $J^{C}_{i,p}$), as well as the corresponding coupling phases ($\phi_{i,p}^{H}$, $\phi_{p}^{V}$, $\phi_{i,p}^{C}$), where $i$ denotes the $i$-th lattice. In the frequency domain, the synthetic lattice points are formed by the resonant modes of each resonator with identical spacing of the free spectral range (FSR) $\Omega$. The phase modulation on the resonators with frequency equal to $p\Omega$ couples the $n$-th and ($n+p$)-th lattice points of each lattice (resonator), signifying the horizontal coupling. The vertical coupling indicates the interaction purely in the spatial domain, i.e., the coupling between the nearest inter-lattice points of the adjacent lattices, which was commonly achieved by a beam splitter with a fixed splitting ratio. To promote convenience and versatility, we replace the splitter with a electro-opitc tunable $2\times 2$ MZI (Fig. 1A). As the splitting ratio of the MZI can be continuously tuned from zero to unity with a direct-current (DC) voltage, the lattices can be arbitrarily adjusted from being completely uncoupled to being fully coupled into a lattice with spacing $\Omega/2$. This advantage can particularly benefit the applications where fine coupling-strength tuning are required, e.g., simulating and extracting Zak phase of SSH model.
@@
Fig. 3: Experimental obtained band structures of the Creutz ladder and direct observation of the Aharonov-Bohm cage effect. (A to E) The heat maps display the general band structures of the Creutz ladder given different $\phi_{1}$ and $\phi_{D}$, where $J^{V}/2J^{H}=1.1$, $J^{C}/J^{H}=0.52$ and $J^{H}=0.06\Omega$. (F) The flat band structure measured at $J^{V}=0$, $J^{H}=J^{C}$, $J^{H}=0.028\Omega$ and $\phi_{A}^{H}=-\phi_{B}^{H}=\pi/2$ when the Aharonov-Bohm cage effect occurs. The lines on the experimental results are the corresponding theoretical bands solved from the $k$-space Hamiltonian. The distribution information on the two spins (resonators) are encoded in the normalized intensities of both the experimental and theoretical normalized intensities, represented by two sets of colors. (G) Illustration of the Aharonov-Bohm cage effect where the phase collected in a round trip in the pink area is $\pi$. When a probe light is input from resonator A with frequency near the zero index mode, the distribution is caged within the dashed box ($0,\pm 1$ in lattice A and ±$\pm 1$ in lattice B). (H and I) The experimental readout of the mode distribution from the drop ports of resonator A and B measured by a Fabry-Perot cavity. The distribution other than those inside the dashed box is clearly suppressed. (J) Illustration of the not caged situation where the phase collected in the pink area is zero. The light normally spreads out along the frequency direction. (K and L) The experimental results of the not caged situation, where more modes survives compared to the caged ones. The distribution results are plotted in the logarithmic coordinates.
@@
Moreover, we highlight that this design makes it possible to deflect the interaction along the spatial direction towards the frequency direction. To be concrete, by adding modulation of frequency $p\Omega$ on the MZI other than the resonators, the mode hops $p$ sites along the frequency direction while crossing to the adjacent resonators, which manifests the cross coupling is satisfied. The introducing of long-range interaction between lattices remarkably raise the complexity of the model and provide opportunity to directly modify the off-diagonal elements of the Hamiltonian matrix in the quasi-momentum space.
Without losing generality, we show experimental demonstration of the fundamental ingredient of the network — the two-resonator protocols in the following. We denote the resonators by A and B, respectively. Retaining the nearest neighbouring coupling in the horizontal and crossing coupling ($p=1$) and setting $\phi^{V}$ to zero, the model turns into the Creutz ladder. We show its schematic in Fig. 1B, of which the Hamiltonian reads
@@
with h.c. representing the Hermitian conjugate.
@@
The experimental setup is illustrated in Fig. 1C. On a TFLN platform, we design and fabricate the double resonators coupled by a MZI with three sets of electrodes. The DC signal applied on the resonators through the Bias-Tees is to align their resonant frequencies. Given the reciprocal relationship between the time domain and the frequency domain, time can be considered as the quasi-momentum ($k$) associated with the frequency lattice. We obtain the $k$-space band structures with the time-resolved spectroscopy approach that is proposed in Ref. [17] exploiting Floquet theory and beyond rotating-wave approximation. Adjusting the detuning $\Delta\omega$ of the probe laser injected into resonator A, we can selectively excite different slices of the band structures. The data is collected from the drop port of the same resonator, which leads to the projection of the band structures on the lattice A. The temporal data for each excited slice allows us to reconstruct the entire projected band structure through data stacking. The intensities of the patterns are proportional to the square of the distribution on lattice A [45]. Additionally, a piezoelectric driven Fabry-Perot cavity is employed as an optical spectrum analyzer to filter specific frequency components within the temporal data, enabling a detailed analysis of the mode distribution.
Tight-binding single lattice and Hall ladder
@@
We start from applying local modulation on the resonators and only DC signals on the MZI ($J^{C}=0$), reducing the Creutz ladder to the Hall ladder. Specially, when the MZI is tuned to disconnect two resonators or to fully couple them, the model further reduces to a tight-binding single lattice.
@@
In Fig. 2A and B, we present the experimentally observed sinusoidal band structures of the tight-binding single lattices. On one hand, when the resonators are not coupled, the lattice constant is $\Omega$ and the nearest-neighbouring lattice points are coupled. On the other hand, when the resonators are fully coupled, the lattice constant is halved and only the next-near-neighbouring lattice points are coupled, doubling the bands. The splitting ratio of the beam splitters in the MZI exhibit a slight deviation from equal division, resulting in imperfections in the experimental outcomes.
@@
Fig. 2E-J illustrates the projected band structures of the Hall ladder with various synthetic magnetic flux $\phi_{1}=2\phi_{A}^{H}=-2\phi_{B}^{H}$ that indicates potential gauges. As the two resonators can be alternatively depicted as pseudospins ($n_{+}$, $n_{-}$), the Hall ladder can signify spin-orbit coupling controlled by the magnetic field. Let $J^{H}=J^{H}_{A}=J^{H}_{B}$ and with the proper choice of $\phi_{1}$, we observe a pronounced dependence of the pseudospin character on $k$. The probe laser selectively excites the band associated with pseudospin $n_{+}$, represented by the incident resonator A (Fig. 2E and F). In the first Brillouin zone, for $k>0$, the excitation occurs in the upper branch, whereas for $k<0$, it occurs in the lower branch. Upon inverting the magnetic flux to $-\phi_{1}$, resonator A turns into pseudospin $n_{-}$ due to symmetry. Consequently, the band structures exhibit mirrored patterns, indicative of chiral spin-momentum locking. Notably, we observe a weakening of this locking effect when the ratio $J^{H}/J^{V}$ decreases as illustrated in Fig. 2I and J.
Creutz ladder
@@
Compared to the two-leg Hall ladder, the full Creutz ladder enables cross coupling between the two lattices, introduces an extra tunable magnetic flux $\phi_{2}$ and $\phi_{3}$, and provides feasibility to combine band flatness and topology. In a similar manner, the cross coupling strength is tuned by the sinusoidal RF signal amplitude and the magnetic flux is modified by the relative phases between three driving signals.
@@
We again set equal coupling strength within each lattice and $\phi_{2}=-\phi_{3}=\phi^{C}-\phi_{A}^{H}$. In Fig. 3A-E, we activate all three types of couplings and alter magnetic flux to experimentally obtain band structures, which are in excellent agreement with the theoretical prediction. To experimentally obtain band flatness, we adjust the DC signal to neutralize the second type of coupling $J^{V}$ and then set $J^{C}=J^{H}$, $\phi_{A}^{H}=-\phi_{B}^{H}=\pi/2$ and $\phi^{C}=0$. The obtained flat band is displayed in Fig. 3F. The Aharonov-Bohm cage effect is one of the direct topological features of the Creutz ladder under the flatness condition, where the wave function is caged within five sites although no boundary on lattices is introduced (Fig. 3G). To directly observe the clues of the Aharonov-Bohm cage effect, in the same situation of Fig. 3F, we inject the laser near the center (zeroth) frequency into resonator A and detect the mode distribution from the drop ports of resonator A and B by a scanning Fabry-Perot cavity. We note that the caging effect is not sensitive to the input laser frequency and we set the laser detuning to approximate $2J^{H}$ for highest signal intensity. More information can be found in Supplementary Materials. Aside from the flat band in the $k$ space, we also observe the cage effect in the real (frequency) space and the experimental results are shown in Fig. 3H and I. The distribution on the zeroth site of lattice B and high order sites (absolute value larger than one) of both lattices are clearly suppressed, featuring the caging effect. As comparison, the experimental results of the not-caged situation where $\phi_{A}^{H}=\phi_{B}^{H}=0$ (Fig. 3J) is displayed in Fig. 3K and L.

### v2

# Versatile photonic frequency synthetic dimensions using a single programmable on-chip device

@@ Abstract
Investigating physical models with photonic synthetic dimensions has been generating great interest in vast fields of science. The rapidly developing thin-film lithium niobate (TFLN) platform, for its numerous advantages including high electro-optic coefficient and scalability, is well compatible with the realization of synthetic dimensions in the frequency together with spatial domain. While coupling resonators with fixed beam splitters is a common experimental approach, it often lacks tunability and limits coupling between adjacent lattices to sites occupying the same frequency domain positions. Here, on the contrary, we conceive the resonator arrays connected by electro-optic tunable Mach-Zehnder interferometers in our configuration instead of fixed beam splitters. By applying bias voltage and RF modulation on the interferometers, our design extends such coupling to long-range scenario and allows for continuous tuning on each coupling strength and synthetic effective magnetic flux. Therefore, our design enriches controllable coupling types that are essential for building programmable lattice networks and significantly increases versatility. As the example, we experimentally fabricate a two-resonator prototype on the TFLN platform, and on this single chip we realize well-known models including tight-binding lattices, the Hall ladder and Creutz ladder. We directly observe the band structures in the quasi-momentum space and important phenomena such as spin-momentum locking, flat band and the Aharonov-Bohm cage effect. These results demonstrate the potential for convenient simulations of more complex models in our configuration.
@@ Introduction
Fig. 1: Schematic and experimental setup. (a) Configuration of a lattice network in frequency synthetic dimensions. The lattice network is simulated by a train of modulated resonators. Instead of fixed beam splitters (BS), MZIs tunable by DC and RF signals are used to couple the adjacent resonators (purple boxes). The lattice network consequently contains three types of coupling $J^{H}_{i,p}$ (orange), $J^{V}_{i}$ (violet), $J^{C}_{i,p}$ (green), where $i$ is the index of resonators ($A_{i}$) along the spatial direction, and $p$ is the coupling range along the frequency direction. These couplings stem from RF signals on the resonators, DC and RF signals applied on the MZIs, respectively. $\phi^{H}_{i,p}$, $\phi^{V}_{i}$, $\phi^{C}_{i,p}$ are the corresponding phases accumulated while coupling, which stem from the initial phases of the modulation. (b) Schematic of the Creutz ladder consisting of two lattices A and B which serve as two pseudospins. The coupling types are of same definition in a and are abbreviated as $J^{H}_{A(B)}$, $J^{V}$ and $J^{C}$. The phases accumulated around the rectangular or triangular plaquettes form gauge potentials such as $\phi_{1}$, $\phi_{2}$ and $\phi_{3}$. (c) Experimental setup for realizing b with two MZI assisted race-track resonators on TFLN. Three sets of electrical signals consisting of DC and RF parts are applied on three sets of the ground-signal-ground (GSG) arranged electrodes through Bias-Tees to control the two resonators ($S_{1},S_{3}$) and the MZI ($S_{2}$). For detecting band structures, a probe laser with detuning $\Delta\omega$ is injected to excite the bands. The time- (quasi-momentum-) varying transmittance signal, which is a slice of the band structure, is detected by a photodetector (PD) followed by an oscilloscope triggered by the RF source. For obtaining mode distribution, the measurement part is replaced by scanning a Fabry-Perot (F-P) cavity together with a photodetector.
@@
Creating a controllable artificial system, of which the dynamics highly resembles the one in real space, is a commonly used method to simulate inaccessible, complex and high-dimensional systems. The artificial systems also called synthetic dimensions, are therefore capable to investigate fascinating novel physics in the field of such as solid-state physics [1, 2, 3, 4]. In terms of realization, photonic systems hold advantages since it has multiple manipulable degrees of freedom including polarization, space, frequency and orbital angular momentum [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Configuring synthetic dimensions in the form of optical frequency lattices is one of the popular approaches in pioneering works, where optical-fiber-based resonators are mainly exploited [17, 18, 19, 20, 21, 16, 22].
@@
Apart from optical fibers, integrated optics holds promise for generating even more complex and programmable lattice networks for its good scalability, stability and reconfigurablity. Especially, with broad transparent spectrum, high non-linear and electro-optic coefficients, thin-film lithium niobate (TFLN) provides an ideal platform. The rapid development of high performance devices on such platform including modulated resonators have paved the way for frequency synthetic dimensions [23, 24, 25, 26, 27, 28, 29, 30, 31]. As a result, constructing frequency lattices on TFLN has garnered increasing interest recently and several experimental works, such as Refs. [32, 33, 34, 35, 36] and our previous work Ref. [37], have been reported.
To advance synthetic dimensions in lattices and interactions, realizing coherent coupling between resonators is crucial. When two resonators are coupled via a passive coupler, adjacent coupling arises between lattice points (resonators) at identical frequencies. This approach facilitates the combination of frequency and spatial dimensions, enabling the simulation of diverse models and intriguing phenomena such as the Hall ladder, the Lieb lattice, band structures with effective magnetic fields, and spin-momentum locking, among others [18, 31, 38]. Furthermore, controllable long-range coupling, particularly cross coupling, is essential for realizing broader classes of models, for example, a historically significant example is the Creutz ladder, one of the earliest models of chiral lattice fermions. This model features edge states protected by spatial inversion symmetry and is characterized by a $\mathbb{Z}$-type topological invariant, even in the absence of chiral symmetry for most cases [39, 40, 41, 42]. Achieving cross coupling requires linking lattice points with distinct frequencies and spatial positions. However, experimental demonstration of such coupling within a frequency-spatial hybrid configuration, employing sufficiently tunable methods, remains an open challenge. Therefore, enhancing the tunability of coupling modes between resonators is necessary to engineer lattice couplings and realize a wider range of physical models.
@@
In this work, we design and fabricate a Mach-Zehnder-interferometer (MZI)-assisted double-resonator device on TFLN, and then realize the sufficient regulation of coupling strength and range in synthetic dimensions. By tuning the bias voltage applied on the MZI [43], we are able to arbitrarily tune the same-frequency coupling strength between two resonators. More importantly, by modulating the MZI rather than resonators with RF signals, we further extend the adjacent coupling to cross coupling and long-range coupling, without introducing auxiliary resonators [11, 44]. This enable us to simulate more complex models like the Creutz ladder. We show the experimental observation of band structures of tight-binding lattices, the Hall ladder and the Creutz ladder, equipped with gauge potentials. Moreover, we also directly observe topological features including the Aharonov-Bohm cage effect which was first carried out by Vidal et al [45, 46, 47, 48]. While maintaining a similar scale and fabrication difficulty to using fixed beam splitters as couplers, our configuration significantly increases the versatility of devices for realizing frequency synthetic dimensions and broadens the scope of models that can be simulated.
Results
Fig. 2: Experimental obtained band structures of tight-binding lattices and the Hall ladder. The band structures exhibit tight-binding lattice characters when two resonators are either not coupled ($J^{V}=0$, a) or fully coupled into a large resonator whose length is doubled and FSR is halved (b). $J_{B}^{H}$ is set to zero in the left panels of (b), while set to be equal to $J_{A}^{H}$ in the right panels. The right panel in (a) and the lower panels in (b) are the corresponding numerical calculations. (c) Illustration of the Hall ladder where the crossing coupling is absent. (d to g) The heat maps are the experimentally obtained band structures of the Hall ladder with coupling strength ratio $r_{VH}=J_{V}/2J_{H}=0.75$, $J^{H}=0.06\Omega$ and an effective magnetic flux $\phi_{1}$ being $\pm\pi/4$ or $\pm 3\pi/4$. (h and i) are the numerical calculations of (f and g). The population information on the two spins (resonators) are encoded in the normalized intensities. Choosing the opposite flux ($\phi_{1}$) is equivalent to selectively exciting the opposite spin. For one spin (d and e), the negative (positive) $k$ states in the first Brillouin zone predominate the upper (lower) bands. Meanwhile, complementary patterns appear in the other spin (f and g), signifying the spin-momentum locking.
@@ Results
Design of the couplings among the lattice network points is one of the core ingredients of the configuration of physical models, specifically, both the coupled modes and coupling range are required to be tuned appropriately. Towards an ideal versatile and programmable network, a device being capable to cover as many coupling types as possible is on demand. Now we consider a 2D network protocol constituted by a series of identical adjacently-coupled 1D lattices, which can be simulated by a train of resonators (Fig. 1a), by integrating the frequency and the spatial domain. We can depict the general model with the horizontal, vertical, crossing coupling strengths ($J^{H}_{i,p}$, $J^{V}_{i}$, $J^{C}_{i,p}$), as well as the corresponding coupling phases ($\phi_{i,p}^{H}$, $\phi_{p}^{V}$, $\phi_{i,p}^{C}$), where $i$ denotes the $i$-th lattice. These phases form effective magnetic fluxes threading the plaquettes of the models, introducing gauge potentials. In the frequency domain, the synthetic lattice points are formed by the resonant modes of each resonator with identical spacing of the free spectral range (FSR) $\Omega$. The phase modulation on the resonators with frequency equal to $p\Omega$ couples the $n$-th and ($n+p$)-th lattice points of each lattice (resonator), signifying the horizontal coupling $J^{H}_{i,p}$. The vertical coupling $J^{V}_{i}$ indicates the interaction purely in the spatial domain, i.e., the coupling between the nearest inter-lattice points of the adjacent lattices, which was commonly achieved by static coupling of two resonators using a beam splitter with a fixed splitting ratio.
To promote convenience and versatility, we replace the splitter with a electro-opitc tunable $2\times 2$ MZI (Fig. 1a) which has two fold of functions. The static coupling is responsible for $J^{V}$. As the splitting ratio of the MZI, standing for static coupling, can be continuously tuned from zero to unity with a direct-current (DC) voltage, the lattices can be arbitrarily adjusted from being statically completely uncoupled ($J^{V}=0$) to being fully coupled into a lattice with spacing $\Omega/2$. This advantage can particularly benefit the applications where fine coupling-strength tuning are required.
@@
Fig. 3: Experimental obtained band structures of the Creutz ladder and direct observation of the Aharonov-Bohm cage effect. (a) Illustration of the Creutz ladder. (b and c) The heat maps display two general band structures of the Creutz ladder given different $\phi_{1}$ and $\phi^{C}$, where $J^{V}/2J^{H}=1.1$, $J^{C}/J^{H}=0.52$ and $J^{H}=0.06\Omega$. (d) The flat band structure measured at $J^{V}=0$, $J^{H}=J^{C}=0.028\Omega$ and $\phi_{A}^{H}=\phi_{B}^{H}=\pi/2$. The absence of $J^{V}$, equal $J^{H}$ and $J^{C}$, together with the $\pi$ flux also satisfy the Aharonov-Bohm cage effect condition. The left panels in (b to d) are the experimental results and the right panels are the numerical calculations. The population information on the two spins (resonators) are encoded in the normalized intensities. (e) Illustration of the Aharonov-Bohm cage effect where the phase collected in a round trip in the pink area is $\pi$. When a probe light is input from resonator A with frequency near the zero index mode, the distribution is caged within the dashed box ($0,\pm 1$ in lattice A and $\pm 1$ in lattice B). (f and g) The blue points are the experimental readout of the mode distribution from the drop ports of resonator A and B measured by a Fabry-Perot (F-P) cavity. The pink dashed lines represent the normalized numerical calculations, where $\Delta\omega$ is fitted to be approximate $4J^{H}$. The distribution other than those inside the dashed box is clearly suppressed. (h) Illustration of the not caged situation where the phase collected in the pink area is zero. The light normally spreads out along the frequency direction. (i and j) The blue points are experimental results of the not caged situation, where more modes survives compared to the caged ones. The pink circles and dashed lines represent the normalized numerical calculations, where $\Delta\omega$ is fitted to be approximate $2J^{H}$. The distribution results are plotted in the logarithmic coordinates and the linewidth of the F-P cavity is taken into account.
@@
Moreover, we highlight that this design makes it possible to deflect the interaction along the spatial direction towards the frequency direction—the modulation of the coupling of two resonators is responsible for $J^{C}_{i,p}$. To be concrete, by adding modulation of frequency $p\Omega$ on the MZI other than the resonators, the mode hops $p$ sites along the frequency direction while crossing to the adjacent resonators, which manifests the cross coupling is satisfied. The introducing of long-range interaction between lattices remarkably raise the complexity of the model and provide opportunity to directly modify the off-diagonal elements of the Hamiltonian matrix in the quasimomentum space.
Without losing generality, we show experimental demonstration of the fundamental ingredient of the network — the two-resonator protocol in the following. We denote the resonators by A and B, respectively. The protocol is capable of simulating 1D tight-binding lattice, as well as quasi-1D two-legged ladder models where the legs (resonators) represent two pseudospins ($_{+}, ss_{-}$), including the Hall ladder and the Creutz ladder. Varying fluxes in ladder systems alters the Hamiltonians, leading to consequences such as time-reversal symmetry breaking, changes of the population information in band structures, interference in hopping processes, and possible potential topological phase transitions. These ladders are one of the simplest models while retaining important properties such as spin-orbit coupling and band flatness. Retaining the nearest neighbouring coupling in the horizontal and crossing coupling ($p=1$) and setting $\phi^{V}$ to zero, the model turns into the Creutz ladder. In the standard Creutz ladder, the parameters are evenly set as $J_{A}^{H}=J_{B}^{H}=J^{C}$, $\phi_{A}^{H}=-\phi_{B}^{H}$ and $\phi^{C}=0$; and also in the standard Hall ladder, $J_{A}^{H}=J_{B}^{H}$ and $\phi_{A}^{H}=-\phi_{B}^{H}$. Here, we still call our models the Creutz ladder or the Hall ladder according to their shapes for simplicity, although our models are more general which can as well include the uneven cases. We show its schematic in Fig. 1b, of which the Hamiltonian in the real space and quasi-momentum ($k$) space reads
@@
$H_{k}=-2J^{H}_{A}\cos(k\Omega+\phi_{A}^{H})\sigma_{+}-2J^{H}_{B}\cos(k\Omega+\phi_{B}^{H})\sigma_{-}$
$-[J^{V}+2J^{C}\cos{(k\Omega+\phi_{C})}]\sigma_{x},$
where h.c. representing the Hermitian conjugate, $\sigma_{x}$, $\sigma_{y}$, $\sigma_{z}$ and $\sigma_{0}$ are Pauli operators on pseudospins, and $\sigma_{\pm}=(\sigma_{0}\pm\sigma_{z})/2$.
@@
The experimental setup is illustrated in Fig. 1c. On a TFLN platform, we design and fabricate the double resonators coupled by a MZI with three sets of electrodes. The DC signal applied on the resonators through the Bias-Tees is to align their resonant frequencies. Given the reciprocal relationship between the time domain and the frequency domain, time can be considered as the quasi-momentum ($k$) associated with the frequency lattice. We obtain the $k$-space band structures with the time-resolved spectroscopy approach that is proposed in Ref. [17]. This can be analyzed utilizing Floquet theory that may be extended beyond the rotating-wave approximation (RWA, but our experiment is typically within the RWA) [17, 18], or scattering matrix approach[49]. The derivation can be found in Supplementary Information. Adjusting the detuning $\Delta\omega$ of the probe laser injected into resonator A, we can selectively excite different slices of the band structures. The data is collected from the drop port of the same resonator, which leads to the projection of the band structures on the lattice A. The time-resolved data for each excited slice enables the reconstruction of the complete projected band structure via data stacking. The pattern intensities are proportional to the square of the population on lattice A. Additionally, a piezoelectrically-driven Fabry-Perot cavity serves as an optical spectrum analyzer to filter specific frequency components from the time-resolved data, enabling detailed analysis of the mode distribution.
@@
Tight-binding single lattice and Hall ladder. We start from applying local modulation on the resonators and only DC signals on the MZI ($J^{C}=0$), reducing the Creutz ladder to the Hall ladder. Specially, when the MZI is tuned to disconnect two resonators or to fully couple them, the model further reduces to a tight-binding single lattice.
@@
In Fig. 2, a and b, we present the experimentally observed sinusoidal band structures of the tight-binding single lattices. On one hand, when the resonators are not coupled, the lattice constant is $\Omega$ and the nearest-neighbouring lattice points are coupled. On the other hand, when the resonators are fully coupled, the effective length of the ring is doubled, halving the FSR (lattice constant). The modulation is kept equal to the original FSR and the next-near-neighbouring lattice points are coupled.
@@
Fig. 2, d-i, illustrates the projected band structures of the Hall ladder with various synthetic magnetic flux $\phi_{1}=-2\phi_{A}^{H}=2\phi_{B}^{H}$ that indicates gauge potentials. As the two resonators can be depicted as pseudospins ($_{+}, ss_{-}$), the Hall ladder can signify spin-orbit coupling controlled by the magnetic field. Let $J^{H}=J^{H}_{A}=J^{H}_{B}$ and with the proper choice of $\phi_{1}$, we observe a pronounced dependence of the pseudospin character on $k$. The probe laser selectively excites the band associated with pseudospin $s_{+}$, represented by the incident resonator A (Fig. 2, d and e). In the first Brillouin zone, for $k>0$, the excitation occurs in the upper branch, whereas for $k<0$, it occurs in the lower branch. Upon inverting the magnetic flux to $\phi_{1}$, resonator A turns into pseudospin $n_{-}$ due to symmetry. Consequently, the band structures exhibit mirrored patterns, indicative of chiral spin-momentum locking.
@@
Creutz ladder. Compared to the two-leg Hall ladder, the full Creutz ladder enables cross coupling between the two lattices, introduces an extra tunable magnetic flux $\phi_{2}$ and $\phi_{3}$, and provides feasibility to combine band flatness and topology. In a similar manner, the cross coupling strength is tuned by the sinusoidal RF signal amplitude and the magnetic flux is modified by the relative phases between three driving signals.
@@
We again set equal coupling strength within each lattice and $\phi_{2}=-\phi_{3}=\phi^{C}-\phi_{A}^{H}$. In Fig. 3, b and c, We activate all three coupling types and vary the magnetic flux to experimentally reconstruct the band structures and show the related numerical calculations. Besides, when parameters are tuned to values at which the topological phase (characterized by the Zak phase) changes, the bandgap undergoes closure. Details regarding the Zak phase of the Creutz ladder and the experimental observation of this gap closing are provided in the Supplementary Information. To experimentally obtain band flatness, we adjust the DC signal to neutralize the second type of coupling $J^{V}$ and then set $J^{C}=J^{H}$, $\phi_{A}^{H}=\phi_{B}^{H}=\pi/2$ and $\phi^{C}=0$. The obtained flat band is displayed in Fig. 3d. The Aharonov-Bohm cage effect is one of the direct topological features of the Creutz ladder under the flatness condition, where the wave function is caged within five sites although no boundary on lattices is introduced (Fig. 3e). We inject laser light near the center frequency (zeroth mode) into resonator A and detect the mode distribution from the drop ports of resonators A and B by scanning the length of the Fabry-Perot cavity. We observe the Aharonov-Bohm cage effect in the real (frequency) space with experimental results shown in Fig. 3, f and g. The distribution exhibits clear suppression at the zeroth site of lattice B and at higher-order sites (absolute mode index larger than 1) on both lattices, characterizing the Aharonov-Bohm cage effect. This effect is insensitive to the input laser frequency and further details are provided in the Supplementary Information. As comparison, the experimental results of the not-caged situation where $\phi_{A}^{H}=\phi_{B}^{H}=0$ (Fig. 3h) is displayed in Fig. 3, i and j. In the numerical calculations, the intensities are normalized according to the experiment.